OpenStudy (anonymous):

Which proportion has a solution of 8? thirty over twelve equals n over forty-five twelve over thirty equals n over forty-five twelve over n equals forty-five over thirty twelve over thirty equals forty-five over n

OpenStudy (anonymous):

okay so im going to translate into numbers

OpenStudy (anonymous):

\[30/12 = n/45 12/30 = n/45\]

OpenStudy (anonymous):

\[12/n = 45/30\] \[12/30 = 45/n\]

OpenStudy (anonymous):

30/12 = n/45 you're gonna want to make n be alone on the side

OpenStudy (anonymous):

that would be 30 * 45/ 12

OpenStudy (anonymous):

112.5, not 8

OpenStudy (anonymous):

lets go straight to 12/n = 45/30

OpenStudy (anonymous):

45/30 = 3/2 = 12/8 n would equal 8

OpenStudy (anonymous):

so your answer would be \[12/n = 45/30\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

not a problem my dear friend